OpenStudy (anonymous):

Log base b 3 = A and log base b 5=B and log base b 7=C What does log base 5b 3 equal in terms of A, B, and C using the change of base fomula What does log base 5b of 35 equal in terms of A,B, and C using the change of base formula

OpenStudy (anonymous):

log base 5b 3=(log base b 3)/(log base b 5b)=A/(log base b 5 +log base b b)=A/(B+1)